# Multi-Derivative Runge-Kutta Flux Reconstruction for hyperbolic conservation laws

### v1

@@ 4.3 Approximate Lax-Wendroff procedure
The time average fluxes $\bm{F}^{\ast}_{p},\bm{F}_{p}$ must be computed using (6), which involves temporal derivatives of the flux. A natural approach is to use the PDE and replace time derivatives with spatial derivatives, but this leads to large amount of algebraic computations since we need to evaluate the flux Jacobian and its higher tensor versions. To avoid this process, we follow the ideas in [42, 6] and adopt an approximate Lax-Wendroff procedure. To present this idea in a concise and efficient form, we introduce the notation
@@
First stage
$\bm{F}^{\ast}:=\bm{f}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{n}\right)=\frac{1}{\Delta t/2}\int_{t^{n}}^{t^{n+1/2}}\bm{f}\left(\bm{u}\right)\text{\text{{{d}}}}t$ (11)
To obtain fourth order accuracy, the approximation for $\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{n}\right)$ needs to be third order accurate (Section A) which we obtain as
@@
Thus, the time averaged flux is computed as
$\mathsf{F}^{\ast}=\mathsf{f}+\frac{1}{4}\mathsf{f}^{(1)}$
@@
$\mathsf{f}^{(1)}$ $=\frac{1}{12}\left[-\bm{f}\left(\mathsf{u}+2\mathsf{u}^{(1)}\right)+8\bm{f}\left(\mathsf{u}+\mathsf{u}^{(1)}\right)-8\bm{f}\left(\mathsf{u}-\mathsf{u}^{(1)}\right)+\bm{f}\left(\mathsf{u}-2\mathsf{u}^{(1)}\right)\right]$
@@
The time averaged flux is computed as
[…]
@@
[…]
@@ 4.4 Numerical flux
$\begin{split}\bm{F}_{e+\frac{1}{2}}&=\frac{1}{2}[\bm{F}_{e+\frac{1}{2}}^{\ast-}+\bm{F}_{e+\frac{1}{2}}^{\ast+}]-\frac{1}{2}\lambda_{e+\frac{1}{2}}[\bm{u}_{h}(x_{e+\frac{1}{2}}^{+},t_{n})-\bm{u}_{h}(x_{e+\frac{1}{2}}^{-},t_{n})]\\ \bm{F}_{e+\frac{1}{2}}&=\frac{1}{2}[\bm{F}_{e+\frac{1}{2}}^{-}+\bm{F}_{e+\frac{1}{2}}^{+}]-\frac{1}{2}\lambda_{e+\frac{1}{2}}[\bm{u}_{h}(x_{e+\frac{1}{2}}^{+},t_{n})-\bm{u}_{h}(x_{e+\frac{1}{2}}^{-},t_{n})]\end{split}$ (12)
@@
$\begin{split}\bm{F}_{e+\frac{1}{2}}^{\ast}&=\frac{1}{2}[\bm{F}_{e+\frac{1}{2}}^{\ast-}+\bm{F}_{e+\frac{1}{2}}^{\ast+}]-\frac{1}{2}\lambda_{e+\frac{1}{2}}[\bm{U}_{e+\frac{1}{2}}^{\ast+}-\bm{U}_{e+\frac{1}{2}}^{\ast-}]\\ \bm{F}_{e+\frac{1}{2}}&=\frac{1}{2}[\bm{F}_{e+\frac{1}{2}}^{-}+\bm{F}_{e+\frac{1}{2}}^{+}]-\frac{1}{2}\lambda_{e+\frac{1}{2}}[\bm{U}_{e+\frac{1}{2}}^{+}-\bm{U}_{e+\frac{1}{2}}^{-}]\end{split}$ (13)
@@
$\begin{split}\bm{U}^{\ast}&=\mathsf{u}+\frac{1}{4}\mathsf{u}^{(1)}\\ \bm{U}&=\mathsf{u}+\frac{1}{6}\left(\mathsf{u}^{(1)}+2\text{{\hskip 1.00006pt}u*{\hskip 1.49994pt}}^{(1)}\right)\end{split}$ (14)
are the time average solutions. In this case, the numerical flux depends on the following quantities: $\{\overline{\bm{u}}_{e}^{n},\overline{\bm{u}}_{e+1}^{n},\bm{U}_{e+\frac{1}{2}}^{-},\bm{U}_{e+\frac{1}{2}}^{+},\bm{F}_{e+\frac{1}{2}}^{-},\bm{F}_{e+\frac{1}{2}}^{+}\}$. Following [3], we will refer to the above two forms of dissipation as D1 and D2, respectively. The dissipation model D2 is not computationally expensive compared to the D1 model since all the quantities required to compute the time average solution $\bm{U}$ are available during the Lax-Wendroff procedure. It remains to explain how to compute $\bm{F}_{e+\frac{1}{2}}^{\pm}$ appearing in the central part of the numerical flux. There were two ways introduced for Lax-Wendroff in [3] to compute the central flux, termed AE and EA. We explain how the two apply to MDRK in the next two subsections.
@@ 4.5 Numerical flux – average and extrapolate to face (AE)
In each element, the time average flux $\bm{F}_{h}^{\delta}$ corresponding to each stage has been constructed using the Lax-Wendroff procedure. The simplest approximation that can be used for $\bm{F}_{e+\frac{1}{2}}^{\pm}$ in the central part of the numerical flux is to extrapolate the flux $\bm{F}_{h}^{\delta}$ to the faces
$\bm{F}_{e+\frac{1}{2}}^{\pm}=\bm{F}_{h}^{\delta}(x_{e+\frac{1}{2}}^{\pm})$
@@ 4.6 Numerical flux – extrapolate to face and average (EA)
Stage 1
$\bm{u}_{\alpha}^{\pm}$ $=\mathsf{V}_{\alpha}^{\top}\left(\mathsf{u}\pm\mathsf{u}^{(1)}\right)$
$\bm{f}^{(1)}_{\alpha}$ $=\frac{1}{12}\left[-\bm{f}\left(\bm{u}^{+2}_{\alpha}\right)+8\bm{f}\left(\bm{u}^{+}_{\alpha}\right)-8\bm{f}\left(\bm{u}^{-}_{\alpha}\right)+\bm{f}\left(\bm{u}^{-2}_{\alpha}\right)\right]$
@@
Stage 2
$\bm{u}_{\alpha}^{\ast\pm}$ $=\mathsf{V}_{\alpha}^{\top}\left(\text{{\hskip 1.00006pt}u*{\hskip 1.49994pt}}\pm\text{{\hskip 1.00006pt}u*{\hskip 1.49994pt}}^{(1)}\right)$
$\bm{u}_{\alpha}^{\ast\pm 2}$ $=\mathsf{V}_{\alpha}^{\top}\left(\text{{\hskip 1.00006pt}u*{\hskip 1.49994pt}}\pm 2\text{{\hskip 1.00006pt}u*{\hskip 1.49994pt}}^{(1)}\right)$
$\bm{f}^{\ast(1)}_{\alpha}$ $=\frac{1}{12}\left[-\bm{f}\left(\bm{u}^{\ast+2}_{\alpha}\right)+8\bm{f}\left(\bm{u}^{\ast+}_{\alpha}\right)-8\bm{f}\left(\bm{u}^{\ast-}_{\alpha}\right)+\bm{f}\left(\bm{u}^{\ast-2}_{\alpha}\right)\right]$
$\bm{F}_{\alpha}$ $=\bm{f}(\bm{u}_{\alpha})+\frac{1}{6}\left(\bm{f}^{(1)}_{\alpha}+2\bm{f}^{\ast(1)}\right)$
@@
The $\bm{f}(\bm{u}_{\alpha})$, $\bm{f}^{(1)}_{\alpha}$ computed in the first stage are reused in the second stage.
@@
$\bm{F}_{e+\frac{1}{2}}^{-}=(\bm{F}_{R})_{e},\qquad\bm{F}_{e+\frac{1}{2}}^{+}=(\bm{F}_{L})_{e+1}$
We will refer to this method with the abbreviation EA. The dissipative part of the flux is computed using either the solution at time $t_{n}$ or the time average solution, which are extrapolated to the faces, leading to the D1 and D2 models, respectively.
@@ 5 Fourier stability analysis
We now perform Fourier stability analysis of the MDRK scheme applied to the linear advection equation $u_{t}+au_{x}=0$ where $a$ is the constant advection speed. We will assume that the advection speed $a$ is positive and denote the CFL number by
@@ 5.1 Stage 1
$\mathsf{u}_{e}^{\ast}=-\sigma\mathsf{A}^{\ast}_{-1}\mathsf{u}_{e-1}^{n}+\left(1-\sigma\mathsf{A}_{0}^{\ast}\right)\mathsf{u}_{e}^{n}-\sigma\mathsf{A}_{+1}^{\ast}\mathsf{u}_{e+1}^{n}$ (16)
@@
$\mathsf{F}_{e}^{\ast}=a\mathsf{U}_{e}^{\ast}\quad\operatorname{where}\quad\mathsf{U}_{e}^{\ast}=\mathsf{T}^{\ast}\mathsf{u}_{e}\quad\operatorname{and}\quad\mathsf{T}^{\ast}=\mathsf{I}-\frac{\sigma}{4}\mathsf{D}$
@@
$F_{h}^{\ast,\delta}(x_{e-\frac{1}{2}}^{+})=\text{{{V}}}_{L}^{T}\mathsf{F}_{e}^{\ast},\qquad F_{h}^{\ast,\delta}(x_{e+\frac{1}{2}}^{+})=\mathsf{V}_{R}^{T}\mathsf{F}_{e}^{\ast}$
@@
$\mathsf{F}_{e+\frac{1}{2}}^{\ast}=\mathsf{V}_{R}^{T}\mathsf{F}_{e}^{\ast}=a\mathsf{V}_{R}^{T}\mathsf{T}^{\ast}\mathsf{u}_{e}$
@@
$F_{e-\frac{1}{2}}^{\ast}-F_{h}^{\ast,\delta}(x_{e-\frac{1}{2}}^{-})=a\mathsf{V}_{R}^{T}\mathsf{u}_{e-1}-a\mathsf{V}_{L}^{T}\mathsf{T}^{\ast}\mathsf{u}_{e},\qquad F_{e+\frac{1}{2}}^{\ast}-F_{h}^{\ast,\delta}(x_{e+\frac{1}{2}}^{-})=0$
@@
$\partial_{\xi}\mathsf{F}_{h}^{\ast}=\mathsf{b}_{L}\left(a\mathsf{V}_{R}^{T}\mathsf{T}^{\ast}\mathsf{u}_{e-1}-a\mathsf{V}_{L}^{T}\mathsf{T}^{\ast}\mathsf{u}_{e}\right)+a\mathsf{D}\mathsf{T}^{\ast}\mathsf{u}_{e}=a\mathsf{b}_{L}\mathsf{V}_{R}^{T}\mathsf{T}^{\ast}\mathsf{u}_{e-1}+a\left(\mathsf{D}\mathsf{T}^{\ast}-\mathsf{b}_{L}\mathsf{V}_{L}^{T}\mathsf{T}^{\ast}\right)\mathsf{u}_{e}$
@@
$\mathsf{u}^{\ast}=\mathsf{u}^{n}-\frac{\Delta t/2}{\Delta x_{e}}\partial_{\xi}\mathsf{F}_{h}^{\ast}$ (17)
@@
$\mathsf{A}_{-1}^{\ast}=\frac{1}{2}\mathsf{b}_{L}\mathsf{V}_{R}^{T}\mathsf{T}^{\ast},\qquad\mathsf{A}_{0}^{\ast}=\frac{1}{2}\left(\mathsf{D}\mathsf{T}^{\ast}-\mathsf{b}_{L}\mathsf{V}_{L}^{T}\mathsf{T}^{\ast}\right),\qquad\mathsf{A}_{+1}^{\ast}=0$ (18)
The upwind character of the D2 dissipation flux leads to $\mathsf{A}_{+1}^{\ast}=0$ and the right cell does not appear in the update equation.

### v2

@@ 4.3 Approximate Lax-Wendroff procedure
The time average fluxes $\bm{F}_{p},\bm{F}^{\ast}_{p}$ must be computed using (7), which involves temporal derivatives of the flux. A natural approach is to use the PDE and replace time derivatives with spatial derivatives, but this leads to large amount of algebraic computations since we need to evaluate the flux Jacobian and its higher tensor versions. To avoid this process, we follow the ideas in [43, 6] and adopt an approximate Lax-Wendroff procedure. To present this idea in a concise and efficient form, we introduce the notation
@@
First stage.
$\bm{F}:=\bm{f}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{n}\right)\approx\frac{1}{\Delta t/2}\int_{t^{n}}^{t^{n+1/2}}\bm{f}\left(\bm{u}\right)\text{\text{{{d}}}}t$ (12)
To obtain fourth order accuracy, the approximation for $\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{n}\right)$ needs to be at least third order accurate (Appendix A) which we obtain as
@@
We denote $\mathsf{f}$ as the vector of flux values at the solution points. Thus, the time averaged flux is computed as
$\mathsf{\bm{F}}=\mathsf{f}+\frac{1}{4}\mathsf{f}^{(1)}$
@@
$\mathsf{f}^{(1)}$ $=\frac{1}{12}\left[-\bm{f}(\mathsf{u}+2\mathsf{u}^{(1)})+8\bm{f}(\mathsf{u}+\mathsf{u}^{(1)})-8\bm{f}(\mathsf{u}-\mathsf{u}^{(1)})+\bm{f}(\mathsf{u}-2\mathsf{u}^{(1)})\right]$
@@
$\bm{F}^{\ast}:=\bm{f}(\bm{u}^{n})+\frac{1}{6}\Delta t\left(\frac{\partial}{\partial t}\bm{f}(\bm{u}^{n})+2\frac{\partial}{\partial t}\bm{f}(\bm{u}^{\ast})\right)$ (13)
The time averaged flux $\bm{F}^{\ast}$ is computed as
$\mathsf{\bm{F}^{\ast}}=\mathsf{f}+\frac{1}{6}(\mathsf{f}^{(1)}+2\mathsf{f}^{*(1)})$
@@
$\mathsf{u}^{*(1)}$ $=-\frac{\Delta t}{\Delta x_{e}}\mathsf{D}\mathsf{f}^{*}$
$\mathsf{f}^{*(1)}$ $=\frac{1}{12}\left[-\bm{f}(\mathsf{u}^{*}+2\mathsf{u}^{*(1)})+8\bm{f}(\mathsf{u}^{*}+\mathsf{u}^{*(1)})-8\bm{f}(\mathsf{u}^{*}-\mathsf{u}^{*(1)})+\bm{f}(\mathsf{u}^{*}-2\mathsf{u}^{*(1)})\right]$
@@ 4.4 Numerical flux
$\begin{split}\bm{F}_{{e+\frac{1}{2}}}&=\frac{1}{2}[\bm{F}_{{e+\frac{1}{2}}}^{-}+\bm{F}_{{e+\frac{1}{2}}}^{+}]-\frac{1}{2}\lambda_{{e+\frac{1}{2}}}[\bm{u}_{h}(x_{{e+\frac{1}{2}}}^{+},t_{n})-\bm{u}_{h}(x_{{e+\frac{1}{2}}}^{-},t_{n})]\\ \bm{F}^{\ast}_{{e+\frac{1}{2}}}&=\frac{1}{2}[\bm{F}^{\ast-}_{{e+\frac{1}{2}}}+\bm{F}^{\ast+}_{{e+\frac{1}{2}}}]-\frac{1}{2}\lambda_{{e+\frac{1}{2}}}[\bm{u}_{h}(x_{{e+\frac{1}{2}}}^{+},t_{n})-\bm{u}_{h}(x_{{e+\frac{1}{2}}}^{-},t_{n})]\end{split}$ (14)
@@
$\begin{split}\bm{F}_{{e+\frac{1}{2}}}&=\frac{1}{2}[\bm{F}_{{e+\frac{1}{2}}}^{-}+\bm{F}_{{e+\frac{1}{2}}}^{+}]-\frac{1}{2}\lambda_{{e+\frac{1}{2}}}[\bm{U}_{{e+\frac{1}{2}}}^{+}-\bm{U}_{{e+\frac{1}{2}}}^{-}]\\ \bm{F}^{\ast}_{{e+\frac{1}{2}}}&=\frac{1}{2}[\bm{F}^{\ast-}_{{e+\frac{1}{2}}}+\bm{F}^{\ast+}_{{e+\frac{1}{2}}}]-\frac{1}{2}\lambda_{{e+\frac{1}{2}}}[\bm{U}^{\ast+}_{{e+\frac{1}{2}}}-\bm{U}^{\ast-}_{{e+\frac{1}{2}}}]\end{split}$ (15)
@@
$\begin{split}\bm{U}^{\ast}&=\mathsf{u}+\frac{1}{4}\mathsf{u}^{(1)}\\ \bm{U}&=\mathsf{u}+\frac{1}{6}\left(\mathsf{u}^{(1)}+2\mathsf{u}^{*(1)}\right)\end{split}$ (16)
are the time average solutions. In this case, the numerical flux depends on the following quantities: $\{\overline{\bm{u}}_{e}^{n},\overline{\bm{u}}_{e+1}^{n},\bm{U}_{{e+\frac{1}{2}}}^{-},\bm{U}_{{e+\frac{1}{2}}}^{+},\bm{F}_{{e+\frac{1}{2}}}^{-},\bm{F}_{{e+\frac{1}{2}}}^{+}\}$. Following [3], we will refer to the above two forms of dissipation as D1 and D2, respectively. The dissipation model D2 is not computationally expensive compared to the D1 model since since all the quantities required to compute the time average solutions $\bm{U},\bm{U}^{\ast}$ are available during the Lax-Wendroff procedure. It remains to explain how to compute $\bm{F}^{\pm}_{e+\frac{1}{2}},\bm{F}^{\ast\pm}_{e+\frac{1}{2}}$ appearing in the central part of the numerical flux. There were two ways introduced for Lax-Wendroff in [3] to compute the central flux, termed AE and EA. We explain how the two apply to MDRK in the next two subsections.
@@ 4.5 Numerical flux – average and extrapolate to face (AE)
In each element, the time average fluxes $\bm{F}_{h}^{\delta}{,\bm{F}^{\ast}_{h}}^{\delta}$ corresponding to each stage have been constructed using the Lax-Wendroff procedure. The simplest approximation that can be used for $\bm{F}^{\pm}_{{e+\frac{1}{2}}},\bm{F}^{\ast\pm}_{{e+\frac{1}{2}}}$ in the central part of the numerical flux is to extrapolate the fluxes $\bm{F}_{h}^{\delta}{,\bm{F}^{\ast}_{h}}^{\delta}$to the faces
$\bm{F}^{\pm}_{{e+\frac{1}{2}}},\bm{F}^{\ast\pm}_{{e+\frac{1}{2}}}=\bm{F}_{h}^{\delta}(x_{{e+\frac{1}{2}}}^{\pm}){,\bm{F}^{\ast}_{h}}^{\delta}(x_{{e+\frac{1}{2}}}^{\pm})$
@@ 4.6 Numerical flux – extrapolate to face and average (EA)
Stage 1.
$\bm{u}_{\alpha}^{\pm}$ $=\mathsf{V}_{\alpha}^{\top}(\mathsf{u}\pm\mathsf{u}^{(1)})$
$\bm{f}^{(1)}_{\alpha}$ $=\frac{1}{12}[-\bm{f}(\bm{u}^{+2}_{\alpha})+8\bm{f}\left(\bm{u}^{+}_{\alpha}\right)-8\bm{f}\left(\bm{u}^{-}_{\alpha}\right)+\bm{f}\left(\bm{u}^{-2}_{\alpha}\right)]$
@@
Stage 2.
$\bm{u}_{\alpha}^{\ast\pm}$ $=\mathsf{V}_{\alpha}^{\top}(\mathsf{u}^{*}\pm\mathsf{u}^{*(1)})$
$\bm{u}_{\alpha}^{\ast\pm 2}$ $=\mathsf{V}_{\alpha}^{\top}(\mathsf{u}^{*}\pm 2\mathsf{u}^{*(1)})$
$\bm{f}^{\ast(1)}_{\alpha}$ $=\frac{1}{12}[-\bm{f}(\bm{u}^{\ast+2}_{\alpha})+8\bm{f}(\bm{u}^{\ast+}_{\alpha})-8\bm{f}(\bm{u}^{\ast-}_{\alpha})+\bm{f}(\bm{u}^{\ast-2}_{\alpha})]$
$\bm{F}_{\alpha}$ $=\bm{f}(\bm{u}_{\alpha})+\frac{1}{6}(\bm{f}^{(1)}_{\alpha}+2\bm{f}^{\ast(1)})$
@@
The $\bm{f}(\bm{u}_{\alpha})$, $\bm{f}^{(1)}_{\alpha}$ computed in the first stage, are reused in the second stage.
@@
$\begin{array}[]{c}\bm{F}^{-}_{e+\frac{1}{2}}=(\bm{F}_{R})_{e},\qquad\bm{F}^{+}_{e+\frac{1}{2}}=(\bm{F}_{L})_{e+1}\\ \bm{F}^{\ast-}_{e+\frac{1}{2}}=(\bm{F}^{\ast}_{R})_{e},\qquad\bm{F}^{\ast+}_{e+\frac{1}{2}}=(\bm{F}^{\ast}_{R})_{e+1}\end{array}$
We will refer to this method with the abbreviation EA.
@@ 5 Fourier stability analysis
We now perform Fourier stability analysis of the MDRK scheme applied to the linear advection equation, $u_{t}+au_{x}=0$, where $a$ is the constant advection speed. We will assume that the advection speed $a$ is positive and denote the CFL number by
@@ 5.1 Stage 1
$\mathsf{u}_{e}^{\ast}=-\sigma\mathsf{A}^{(1)}_{-1}\mathsf{u}_{e-1}^{n}+(1-\sigma\mathsf{A}^{(1)}_{0})\mathsf{u}_{e}^{n}-\sigma\mathsf{A}^{(1)}_{+1}\mathsf{u}_{e+1}^{n}$ (18)
@@
$\mathsf{\bm{F}}_{e}=a\mathsf{\bm{U}}_{e}\quad\operatorname{where}\quad\mathsf{\bm{U}}_{e}=\mathsf{T}^{(1)}\mathsf{u}_{e}\quad\operatorname{and}\quad\mathsf{T}^{(1)}=\mathsf{I}-\frac{\sigma}{4}\mathsf{D}$
@@
$F_{h}^{\delta}(x_{{e-\frac{1}{2}}}^{+})=\text{{{V}}}_{L}^{T}\mathsf{\bm{F}}_{e},\qquad F_{h}^{\delta}(x_{{e+\frac{1}{2}}}^{+})=\mathsf{V}_{R}^{T}\mathsf{F}_{e}$
@@
$\mathsf{\bm{F}}_{{e+\frac{1}{2}}}=\mathsf{V}_{R}^{T}\mathsf{\bm{F}}_{e}=a\mathsf{V}_{R}^{T}\mathsf{T}^{(1)}\mathsf{u}_{e}$
@@
$F_{{e-\frac{1}{2}}}-F_{h}^{\delta}(x_{{e-\frac{1}{2}}}^{+})=a\mathsf{V}_{R}^{T}\mathsf{u}_{e-1}-a\mathsf{V}_{L}^{T}\mathsf{T}^{(1)}\mathsf{u}_{e},\qquad F_{{e+\frac{1}{2}}}-F_{h}^{\delta}(x_{{e+\frac{1}{2}}}^{-})=0$
@@
$\partial_{\xi}\mathsf{\bm{F}}_{h}=\mathsf{b}_{L}(a\mathsf{V}_{R}^{T}\mathsf{T}^{(1)}\mathsf{u}_{e-1}-a\mathsf{V}_{L}^{T}\mathsf{T}^{(1)}\mathsf{u}_{e})+a\mathsf{D}\mathsf{T}^{(1)}\mathsf{u}_{e}=a\mathsf{b}_{L}\mathsf{V}_{R}^{T}\mathsf{T}^{(1)}\mathsf{u}_{e-1}+a(\mathsf{D}\mathsf{T}^{(1)}-\mathsf{b}_{L}\mathsf{V}_{L}^{T}\mathsf{T}^{(1)})\mathsf{u}_{e}$
@@
$\mathsf{u}^{\ast}=\mathsf{u}^{n}-\frac{\Delta t/2}{\Delta x_{e}}\partial_{\xi}\mathsf{\bm{F}}_{h}$ (19)
@@
$\mathsf{A}^{(1)}_{-1}=\frac{1}{2}\mathsf{b}_{L}\mathsf{V}_{R}^{T}\mathsf{T}^{(1)},\qquad\mathsf{A}^{(1)}_{0}=\frac{1}{2}(\mathsf{D}\mathsf{T}^{(1)}-\mathsf{b}_{L}\mathsf{V}_{L}^{T}\mathsf{T}^{(1)}),\qquad\mathsf{A}^{(1)}_{+1}=0$ (20)
The upwind character of the D2 dissipation flux leads to $\mathsf{A}^{(1)}_{+1}=0$ and the right cell does not appear in the update equation.